# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v1

@@ 9. Auxiliary Lemmas
$|a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}$
and $|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}.$
Since $\|\|\bm{w}_{r}(0)\|_{2}^{3}\|_{\psi_{\frac{2}{3}}}\leq(\|\|\bm{w}_{r}(0)\|_{2}^{2}\|_{\psi_{1}})^{\frac{3}{2}}\lesssim d^{3}$, from Lemma 8, we have that for fixed $i\in[n_{1}]$ and $j\in[n_{2}]$ with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}$
@@
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}.$
Then taking a union bound for all $i\in[n_{1}]$ and $j\in[n_{2}]$ with $2(n_{1}+n_{2})e^{-t}=\delta$ yields that
$L(0)$ $\lesssim(d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}})^{2}$
$\lesssim d^{6}t+\frac{d^{6}t^{3}}{m}$
$=d^{6}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{1}{m}\log^{3}\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
$\lesssim d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
as $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$.

### v3

@@ 11. Auxiliary Lemmas
$\left|a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim\|\bm{w}_{r}(0)\|_{2}^{2}|\bm{w}_{r}(0)^{T}\bm{x}_{p}|$
and $\left|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim\|\bm{w}_{r}(0)\|_{2}^{2}|\bm{w}_{r}(0)^{T}\bm{y}_{j}|.$
Since $\left\|\|\bm{w}_{r}(0)\|_{2}^{2}\right\|_{\psi_{1}}=\mathcal{O}(d)$ and $\|\bm{w}_{r}(0)^{T}\bm{y}_{j}\|_{\psi_{2}},\|\bm{w}_{r}(0)^{T}\bm{x}_{p}\|_{\psi_{2}}=\mathcal{O}(1)$, from Lemma D.2, we have that
$\|\|\bm{w}_{r}(0)\|_{2}^{2}|\bm{w}_{r}(0)^{T}\bm{x}_{p}|\|_{\psi_{\frac{2}{3}}}=\mathcal{O}(d),|\bm{w}_{r}(0)^{T}\bm{y}_{j}|\|_{\psi_{\frac{2}{3}}}=\mathcal{O}(d).$
Applying Lemma D.1 yields that for fixed $p\in[n_{1}]$ and $j\in[n_{2}]$ with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim d\sqrt{t}+\frac{d}{\sqrt{m}}t^{\frac{3}{2}}$
@@
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim d\sqrt{t}+\frac{d}{\sqrt{m}}t^{\frac{3}{2}}.$
Then taking a union bound for all $p\in[n_{1}]$ and $j\in[n_{2}]$ with $2(n_{1}+n_{2})e^{-t}=\delta$ yields that
$L(0)$ $\lesssim\left(d\sqrt{t}+\frac{d}{\sqrt{m}}t^{\frac{3}{2}}\right)^{2}$
$\lesssim d^{2}t+\frac{d^{2}t^{3}}{m}$
$=d^{2}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{1}{m}\log^{3}\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
$\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
since $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$.